OpenStudy (anonymous):

In 5 years, Dad will be three times as old as his daughter Jill will be then. If the sum of their present ages is 50, how old are they now? If x + 5 is Jill's age five years from now, which of the following equations could be used to solve the problem? 3x + 5 = 55 - x 3x + 15 = 55 - x 3x + 15 = 45 - x

OpenStudy (anonymous):

let D = dad's age now. let J = jill's age now. D+5 = dad's age 5 years from now. J+5 = jill's age 5 years from now. problem states that the sum of their ages now is 50, so D+J=50 problem states that in 5 years dad will be 3 times as old as jill will be then, so (D+5) = 3 * (J+5) to solve this problem, we remove one of the unknowns by expressing it in terms of the other unknown in one of the equations. i took D+J=50 because it was simpler. i could have used the other equation as well. either way it will work. if D+J=50, then D=(50-J) taking that relationship and substituting for D in the other equation, it becomes 50 - J + 5 = 3 * (J + 5) this becomes 50 - J + 5 = (3 * J) + (3 * 5)

OpenStudy (anonymous):

DId that help?

OpenStudy (anonymous):

yes it did thank you